# Leave No Patient Behind: Enhancing Medication Recommendation for Rare Disease Patients

### v1

@@ B.1. Baseline Details
B.2. Metrics
In this subsection, we define the metrics employed within the experiment: Jaccard Similarity Score (Jaccard), Precision Recall AUC (PRAUC), F1 score (F1), and DDI rate (DDI). Building upon the formulation detailed in Section 3, $\mathcal{M}$ represents the set of all medications in the dataset, while $\mathbf{m}^{(j)},\hat{\mathbf{m}}^{(j)}\in\{0,1\}^{|\mathcal{M}|}$ signify the prescribed and predicted medications to a patient identified by index $j$, respectively. For simplicity, we will subsequently omit the superscript $j$ for a single-visit patient.
  • Jaccard:
  • F1:
    $\text{F}_{1}=\frac{2\text{R}\times\text{P}}{\text{R}+\text{P}},$
    where the recall and precision are formulated as
  • PRAUC:
  • DDI:
    where $\mathbf{A}$ is the DDI graph defined in Section 3.
Table 5. Medication recommendation results for a rare disease patient. Here, ”FN” (False Negative) denotes medications present in the ground truth but not predicted, whereas ”FP” (False Positive) refers to predicted medications not found in the ground truth. The underlined diseases appear less than 30 times in the experimental dataset.
Diagnoses & Procedures (ICD codes (Slee, 1978)) Methods Medication Recommendation (ATC codes (Skrbo et al., 1999))
Ground Truth A02A, N02B, B05C, A12A, A12C, A06A, B01A, A04A, N06A, N05C, B03B, N05A, A02B
Diagnoses: 5799, 71943, 5959, 920, 61610, J01D, N05B, J01E, N06D, A11D, H04A, A12B, N03A, A01A (totally 22 medications)
30392, 80122, 37941, 8020, 8054, 2811, 311, GAMENet 14 correct, 8 FN (A02A, A11D, A12A, B03B, H04A, J01E, N05C, N06D), 2 FP (A07A, N02A)
E8844, 29420, V1271, 04149, E8809, E8796, SafeDrug 16 correct, 6 FN (A02A, A01A, A11D, J01E, N05C, N06D), 3 FP (C07A, N02A, N07B)
27650, V1588, 99664, 29181, 34982, 78791, MICRON 17 correct, 5 FN (A02A, J01E, N03A, N05C, N06D), 3 FP (A07A, C07A, N02A)
42789, 2875, 2762, 2809, 2753, 2930 (totally COGNet 17 correct, 5 FN (A02A, A12A, J01E, N05C, N06D), 5 FP (A07A, C07A, C10A, R01A, R03A)
30 diseases) MoleRec 17 correct, 5 FN (A02A, J01E, N05C, N06A, N06D), 4 FP (A07A, C07A, N02A, N07B)
Procedures: 8659, 3897 $w/o$ P 17 correct, 5 FN (A02A, A11D, J01E, N05C, N06D), 5 FP (A07A, C07A, N02A, N07B, R03A)
RAREMed 19 correct, 3 FN (J01E, N05C, N06D), 2 FP (A07A, N02A)
Appendix C Exploratory Analysis
C.1. Case Study
To demonstrate RAREMed’s effectiveness in handling rare cases, we select a patient from the rare group in the test set of MIMIC-III, whose characteristics are detailed in Table 5, revealing the presence of 30 diagnosed diseases, 2 procedures, and 22 prescribed medications. Notably, among the diagnosed diseases, 11 of them appear less than 30 times within the dataset, indicating their rarity. We exclude G-Bert (Shang et al., 2019a) and RETAIN (Choi et al., 2016) as they cannot deal with such single-visit patient.
Remarkably, RAREMed outperforms all baselines and the ”$w/o$ P” variant, achieving 19 accurate predictions. It encounters only 3 medication oversights and makes 2 incorrect predictions, thereby showcasing its accurate performance in this challenging scenario.
Specifically, it is noteworthy that all models, except RAREMed, fail to identify the medication A02A, which corresponds to antacids used to alleviate stomach pain and digestive issues, and holds strong relevance with the disease ”unspecified digestive system disorder” (indicated by the ICD code 5799).
These findings underscore the superiority of RAREMed in providing precise medication recommendations for patients with rare diseases, thereby showcasing its potential in addressing the fairness concern within the realm of medication recommendation.
C.2. Importance of Single-Visit Patients
In this section, we present the rationale behind our choice to develop RAREMed as an instance-based method, focusing on the effective utilization of clinical information from the current visit. We discuss this decision from two key perspectives:
  • Dominance of single-visit patients in EHR data: Real EHR datasets exhibit a prominent trend, as depicted in Figure 6 (a), wherein a significant majority of patients have only a single visit recorded. This observation accurately reflects the prevailing real-world scenario, where most patients are first-time visitors. Moreover, privacy considerations further hinder the comprehensive utilization of patients’ historical information. Therefore, it is crucial to develop methods that can effectively leverage the clinical information available at the current visit to provide accurate and reliable recommendations for this majority population.
  • Challenges in leveraging historical information: Existing approaches that rely on historical information often fail to deliver satisfactory results due to incomplete exploitation of such data. Notably, models such as GAMENet (Shang et al., 2019b), SafeDrug (Yang et al., 2021b) and MoleRec (Yang et al., 2023) leverage RNN to encode historical information, while the accuracy does not exhibit significant growth as the number of visits increases, as shown in Figure 6(b). COGNet (Wu et al., 2022) introduces a copy-or-predict mechanism to copy medication information from similar historical visits. However, this approach can be easily replicated by doctors, thereby diminishing its value. MICRON (Yang et al., 2021a) focuses on medication changes between visits, but the prediction errors tend to accumulate as the number of visits increases, evidenced by the decreasing trend in Figure 6(b). In summary, effectively leveraging historical information remains a challenging task, which we acknowledge as an important area for future research.
In summary, our focus on single-visit patients in developing RAREMed is justified by the dominance of such patients in real-world EHR data and the challenges associated with effectively utilizing historical information. By prioritizing the clinical information available at the current visit, we aim to provide accurate and reliable medication recommendations that address the specific needs of this majority population.
Figure 6. (a) Skewed distribution of #visits in EHR. (b) Medication accuracy across different visit numbers of several representative longitudinal-based methods on MIMIC-III.

### v2

@@ 6.1. Experimental Setup
In addition to the medication recommendation methods mentioned above, we also implement a baseline approach specifically designed for fair recommendation in group-wise comparison:
  • Rebalancing: We employ a resampling technique (Kamiran and Calders, 2009; Ekstrand et al., 2018) to address the issue of data imbalance. Specifically, we calculate the Inverse Propensity Score (IPS) score (Schnabel et al., 2016; Yang et al., 2018) for each disease code. Subsequently, we assign each patient a weight score based on the IPS score associated with the rarest disease. In each training epoch, we perform random sampling with replacement from the training data, the probability of selecting a particular patient is proportional to their assigned patient weight. We conduct experiment for this approach on our RAREMed without pre-training.
Implementation Details. The hyper-parameters of all models are determined based on their performance on the validation set. RAREMed is pre-trained on the training set in a sequential manner, first on the SMP task and then on the SP task, for 30 epochs respectively. The transformer encoder in our model consists of 3 layers with 4 attention heads, and the embedding dimension is set to 512. The weights of the loss function, denoted as $\alpha$ and $\beta$, are set to 0.03 and 0.7, respectively. The parameters are trained using the AdamW optimizer (Loshchilov and Hutter, 2017) with a learning rate of 1e-5 and weight decay of 0.1.
6.2. Overall Performance Comparison (RQ1)
Model Overall Performance Group-wise Performance
$\mathbf{w/o}$ P 0.4561 0.6939 0.6092 0.0509 12.94 0.4719 0.4666 0.4534 0.4475 0.4410 0.01157
$\mathbf{w/o}$ U 0.4589 0.6965 0.6122 0.0545 13.13 0.4773 0.4707 0.4558 0.4485 0.4423 0.01320
$\mathbf{w/o}$ s&r 0.4528 0.6892 0.6062 0.0554 13.15 0.4695 0.4646 0.4508 0.4451 0.4342 0.01286
  • Baseline Comparison. LR and RETAIN, which were not originally designed for medication recommendation tasks, exhibit poor performance. LR overlooks the interaction among clinical codes, while RETAIN fails to adequately capture the current patient condition. Similarly, LEAP and G-Bert, which do not consider the procedure sequence, also underperform in this domain. Furthermore, COGNet and MICRON utilize historical information in a specific manner, but their performance significantly declines in the single-visit setting. By integrating external knowledge such as EHR graph and drug molecule structures, GAMENet, SafeDrug and MoleRec achieve better performance. Interestingly, these three methods perform even worse in multi-visit setting compared to single-visit setting with MIMIC-III, highlighting the complexity of leveraging historical information.
6.3. Group-wise Performance Comparison (RQ2)
  • Baseline Methods Yield Unfair Recommendations. Conventional medication recommendation methods exhibit a notable decrease in accuracy as the frequency of disease codes decreases, as illustrated in Figure 4(a). This leads to a higher standard deviation of Jaccard across different groups, as depicted in Figure 4(b). Notably, G-Bert (Shang et al., 2019a), which utilizes pre-training techniques, struggles to achieve satisfactory accuracy for patients with rare diseases, potentially due to its disregard for procedure codes and subsequent failure to capture a comprehensive representation of patients’ clinical information during pre-training. This underscores the superiority of our model structure and pre-training task design. Additionally, Rebalancing, tailored to address fairness concerns, also underperforms compared to RAREMed in terms of both fairness and overall performance.
  • Pre-training Demonstrates Effectiveness on Biased Datasets. The results presented in Figure 4(c) and (d) highlight the overall positive impact of our pre-training strategy on both common group and rare group, with few exceptions. Specifically, on MIMIC-III, where the disease distribution is less skewed, our pre-training strategy consistently enhances performance across various groups, as evidenced by the uniform improvement ratio observed after 30 pre-training epochs. In contrast, on MIMIC-IV, characterized by a more biased disease distribution and presenting greater challenges666 This is supported by the extensive distribution of long-tail disease codes in MIMIC-IV, as illustrated in Figure 1, and the greater variance of Jaccard scores across groups, as presented in Table 4, in comparison to the findings in Figure 4(b). , RAREMed notably improves the performance of the rarest group, effectively mitigating unfairness. The substantial improvement in overall accuracy on such a more biased dataset is also noteworthy.
6.4. Ablation Study (RQ3)
  • $\mathbf{w/o}$ P”, which removes two pre-training tasks, resulting in initializing parameters of encoder and embedding layers randomly.
  • $\mathbf{w/o}$ U”, which disables the unified encoder and learns two encoders for diseases and procedures seperately and then concatenates the output. This setup retains two pre-training tasks.
  • $\mathbf{w/o}$ s&r”, it disables segment and relevance embedding layers, and patient representation relies solely on the token embedding.
  • $\mathbf{w/o}$ P”. As anticipated, the diminished performance observed in the ”w/o P” variant underscores the importance of our pre-training tasks. Notably, RAREMed outperforms ”$w/o$ P” across all five groups, affirming that our pre-training strategy not only enhances performance on rare diseases but also preserves the efficacy of the original model on other diseases. The enhancement seen in RAREMed can be attributed to the augmented encoder obtained through the pre-training phase, enabling RAREMed to grasp intricate disease/procedure associations and produce more comprehensive patient representations.
6.5. Hyperparameter Studies (RQ4)
Loss Weight. Figure 5(a) demonstrates that the value of $\alpha$ significantly affects the size of the recommended medicine set. Additionally, we observed that the recommendation accuracy initially increases and then declines as the number of recommended medications increases. This suggests that in addition to the ranking task commonly considered in classic recommendation systems, the size of recommended medication set is also an important factor in the field of medication recommendation.
Number of Pre-training Epochs. Figure 5(c) illustrates the impact of the number of pre-training epochs on the pre-training tasks and the medication recommendation task. It is evident that as the number of pre-training epochs increases, the model’s performance in pre-training tasks (SR task and SMP task) improves. This observation indicates that the model acquires knowledge about clinical code associations and gains comprehensive patient representations during pre-training, which are crucial for medication recommendation, particularly for patients with rare diseases. However, the improvement in the downstream task (Jaccard) starts to diminish after approximately 30 epochs. This phenomenon may be attributed to overfitting in the pre-training task and the inherent differences between the pre-training and downstream tasks. These findings highlight the significance of striking an optimal balance between the performance on pre-training tasks and the downstream medication recommendation task.
Figure 5. Hyperparameter effects on model performance.
7. Conclusion
In this paper, we propose RAREMed, a novel medication recommendation model designed to addresses the fairness issue in medication recommendation systems. By focusing on enhancing accuracy for patients with rare diseases, RAREMed leverages two self-supervised pre-training tasks to learn specialized medication needs and interrelations. The model also utilizes a unified input sequence approach to capture complex relationships among disease and procedure codes. Experimental results on real-world datasets demonstrate the effectiveness of RAREMed in providing accurate drug recommendations for both rare and common disease patients, thereby mitigating unfairness.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (62272437, 62302321) and the University Synergy Innovation Program of Anhui Province (GXXT-2023-071).